OpenStudy (sabasattar):

A new drug is designed to reduce a person's blood pressure. Eight randomly selected hypersensitive patients receive the new drug. Suppose the probability that the hypersensitive patient's blood pressure drops if he/she receive the new drug is 0.5. What is the mean and standard deviation of this distribution?

OpenStudy (518nad):

what is the defn of STD again

OpenStudy (sabasattar):

STD?

OpenStudy (518nad):

standard deviation

OpenStudy (sabasattar):

"a quantity calculated to indicate the extent of deviation for a group as a whole."

OpenStudy (518nad):

i want the formula for it

OpenStudy (sabasattar):

That's what I am confused about. Which formula should I use? There's several

OpenStudy (518nad):

theres one main formula for STD

OpenStudy (sabasattar):

Could you please share what it is?

OpenStudy (sabasattar):

Thank you. Is there an online calculator I could use?

OpenStudy (518nad):

your sample size is pretty small, so ur mean is 4

OpenStudy (sabasattar):

Most STD calculators require the range of data set but this one doesn't have it

OpenStudy (sabasattar):

Is the STD 2?

OpenStudy (518nad):

u can generate it

OpenStudy (sabasattar):

I see! Thank you so much!

OpenStudy (518nad):

umm lemme see

OpenStudy (sabasattar):

I have one last question!

OpenStudy (518nad):

sqrt(npq) = sqrt(8*1/2*1/2) = sqrt(4) = 2

OpenStudy (518nad):

ya it is 2

OpenStudy (518nad):

i cant tell u why that formula works exactly, but there is a way to work it out

OpenStudy (518nad):

but when ur given a sample size, and its one of these success failure ones then sqrt(npq) is the STD

OpenStudy (sabasattar):

The following data represent a random sample of 9 scores on a statistics quiz: Quiz Scores: 7,9,7,5,5,8,3,10,9 The scores are known to have a Normal distribution and the sample standard deviation is s=2.29. Construct a 95% confidence interval to estimate the population mean.

OpenStudy (518nad):

where n is sample size,p is success and q is failure prob

OpenStudy (518nad):

okay

OpenStudy (518nad):

you need to know the relationship between sample SD and true SD